PROBLEM SOLVING: hOW TO SoLve
Suppose your class recieves a $1084 for seeling 205 packages of greeting cards and gift wrap. If tyhe girft wrap is 4 dollars a package and the greeting cards are 10 a package, find the number of each type of package sold.

Answer by checkley71(8403)

C+W=205 OR C=205-W
4C+10W=1084
4(205-W)+10W=1084
820-4W+10W=1084
6W=1084-820
6W=264
W=264/6
W=44 NUMBER OF WRAP ITEMS.
C+44=205
C=205-44
C=161 NUMBER OF CARDS SOLD.
PROOF
4*161+10*44=1084
644+440=1084
1084=1084